OpenStudy (anonymous):

x - 2y = 3 5x + 3y = 2 The lines whose equations are shown intersect at (1, -1) (-1, 1) (0, 3/2)

OpenStudy (igreen):

x - 2y = 3 Add 2y to both sides: x = 2y + 3 Plug in '2y + 3' for 'x' in the 2nd equation: 5(2y + 3) + 3y = 2 Distribute 5 into the parenthesis: 10y + 15 + 3y = 2 Combine like terms, what's 10y + 3y?

OpenStudy (ribhu):

i hope u got it!